# Spurious reconstruction from brain activity

### v1

@@ 4.6 Expected identification accuracy in imprecise reconstructions
A pairwise identification accuracy is a metric defined on three types of samples: the test sample, the predicted sample, and the candidate sample selected from a test set as $\mathbf{y},\mathbf{\hat{y}},\mathbf{y}_{-}\in\mathcal{Y}$, respectively. We define a function $S\colon\mathcal{Y}\times\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$ that takes the aforementioned triplet as the input and output whether the predicted sample was much close to the test sample than the candidate sample as
@@
$\mathrm{Acc}=\frac{1}{n(n-1)}\sum_{i}^{n}\sum_{j\neq i}^{n}S(\hat{\mathbf{y}}^{(i)},\mathbf{y}^{(i)},\mathbf{y}_{-}^{(i)}).$ (12)
Now we consider a scenario where the translator only decodes semantic information (e.g., category) and cannot decode information about its precise visual appearance. Suppose the test set contains a categorical structure like NSD stimuli, we model such a scenario as
@@
$Z$ is a set of triplet of which the test sample and the candidate sample belong to the same category. $\bar{Z}$ is a complementary set of $Z$. $\mathbb{E}_{p(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-}\mid\cdot)}[S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})]$ represents the pairwise identification accuracy in the conditional expectation form. If the candidate sample belongs to the same category as the test sample, pairwise identification is challenging because of the poor prediction of the translator. On the other hand, if the candidate samples belongs to a different category than the test sample, the test sample is easily identified only from the semantic information.
Here, we assume that the test set contains $k$ categories in total and all samples are equally distributed across each category for simplicity. If we have sufficiently large number of test samples, the above identification accuracy can be approximated as
$\mathrm{Acc}$ $\underset{\phantom{n\to\infty}}{=}\frac{1}{n(n-1)}\left(\sum_{(\mathbf{y},\mathbf{\hat{y}},\mathbf{y}_{-})\in Z}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)+\frac{1}{n(n-1)}\left(\sum_{(\mathbf{y},\mathbf{\hat{y}},\mathbf{y}_{-})\in\bar{Z}}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)$ (15)
$\underset{\phantom{n\to\infty}}{=}\frac{|Z|}{n(n-1)}\left(\frac{1}{|Z|}\sum_{(\mathbf{y},\mathbf{\hat{y}},\mathbf{y}_{-})\in Z}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)+\frac{|\bar{Z}|}{n(n-1)}\left(\frac{1}{|\bar{Z}|}\sum_{(\mathbf{y},\mathbf{\hat{y}},\mathbf{y}_{-})\in\bar{Z}}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)$ (16)
=$\underset{n\to\infty}{=}\frac{1}{k}\cdot\mathbb{E}_{p(\mathbf{\hat{y}},\mathbf{y},\mathbf{y}_{-}\mid(\mathbf{y},\mathbf{\hat{y}},\mathbf{y}_{-})\in Z)}[S]+\left(1-\frac{1}{k}\right)\cdot\mathbb{E}_{p(\mathbf{\hat{y}},\mathbf{y},\mathbf{y}_{-}\mid(\mathbf{y},\mathbf{\hat{y}},\mathbf{y}_{-})\in\bar{Z})}[S]$ (17)
@@
where $|Z|=n\left(n/k-1\right)$, and $|\bar{Z}|=n\left(n-n/k\right)$. We used the assumption on the large sample size at the third equality.
5 Acknowledgements
@@
We thank our laboratory team, especially Eizaburo Doi, Hideki Izumi, Matthias Mildenberger and for their invaluable comments and suggestions on the manuscript. This work was supported by the Japan Society for the Promotion of Science (JSPS: KAKENHI grants JP20H05954, JP20H05705 to Y.K, JP21K17821 to Y.N, and 22KJ1801 to K.S.), Japan Science and Technology Agency (JST: CREST grants JPMJCR18A5, and JPMJCR22P3 to Y.K.), and New Energy and Industrial Technology Development Oprganization (NEDO: commissioned project, JPNP20006 to Y.K.).
Appendix A Supplementary figures
Figure S1: UMAP visualization of CLIP text features in Deeprecon. Blue points represent training samples, while red points represent test samples.
Figure S2: Supplementary information of hold out analysis. (a) The clustering results, applying the $k$-Means clustering to the UMAP embedding of CLIP text features. Based on these clustering results, the assignment of training and test samples under the hold-out split condition was determined. (b) Confusion probability matrix of cluster identification results. The top row shows the cluster identification results under the hold-out split conditions. The left side is for CLIP text features and the right side is for CLIP vision features. The bottom row shows the cluster identification results under the naive split conditions. In each matrix, the $(i,j)$ cell indicates the ratio of samples assigned by cluster $i$ are predicted to cluster $j$. The empty cell means the ratio is 00.
Figure S3: Additional results about simulation with clustered data. (a) Cluster identification accuracy as a function of the cluster variance ratio. The $x$ axis indicates the cluster variance ratio $\sigma_{r}^{2}/\sigma_{s}^{2}$ and the $y$ axis represents the cluster identification performance. Green lines show the results of in-distribution test samples and orange lines show those for OOD samples. The number of clusters used in training is increasing as the color lightens. Similar to Figure 9b in the main figure, the OOD identification performance become closer to the in distribution samples as the cluster variance ratio is increasing. (b). Heatmap plot of cluster identification accuracy varying the number of clusters in training and dimension. The left matrix shows the cluster identification results for the in-distribution test samples. The right matrix shows the cluster identification results for the OOD test samples. In each matrix, the $(i.j)$ cell indicates the cluster identification accuracy when the dimension is $i$ and the number of clusters used in the training is $j$.
@@ Appendix A Supplementary figures
Figure S4: Frame simirality in Nishimoto et al. (2011). We first detected scenes from the training and test movie stimuli. Then, for the first and last frames of each test scene, we searched for the frame from the training movies with the closest Euclidean distance. Our analysis revealed that 37 out of the 48 scenes in the test set contained frames that were nearly identical to those in the training set.

### v3

@@ Discussion
A pairwise identification accuracy is a metric defined on three types of samples: the test sample, the predicted sample, and the candidate sample selected from a test set as $\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-}\in\mathcal{Y}$, respectively. We define a function $S\colon\mathcal{Y}\times\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$ that takes the triplet above as the input and output whether the predicted sample was much closer to the test sample than the candidate sample as
@@
$\mathrm{Acc}=\frac{1}{n(n-1)}\sum_{i}^{n}\sum_{j\neq i}^{n}S(\hat{\mathbf{y}}^{(i)},\mathbf{y}^{(i)},\mathbf{y}_{-}^{(j)}).$ (10)
Now, we consider a scenario where the translator only decodes semantic information (e.g., category) and cannot decode information about its precise visual appearance. Suppose the test set contains a categorical structure like NSD stimuli, we model such a scenario as
@@
$Z$ is a set of triplets in which the test sample and the candidate sample belong to the same category. $\bar{Z}$ is a complementary set of $Z$. $\mathbb{E}_{p(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-}\mid\cdot)}[S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})]$ represents the pairwise identification accuracy in the conditional expectation form. If the candidate sample belongs to the same category as the test sample, pairwise identification is challenging because of the poor prediction of the translator. On the other hand, if the candidate samples belong to a different category than the test sample, the test sample is easily identified only from the semantic information.
Here, we assume that the test set contains $k$ categories in total and that all samples are equally distributed across each category for simplicity. If we have a sufficiently large number of test samples, the above identification accuracy can be approximated as
$\mathrm{Acc}$ $\underset{\phantom{n\to\infty}}{=}\frac{1}{n(n-1)}\left(\sum_{(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in Z}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)+\frac{1}{n(n-1)}\left(\sum_{(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in\bar{Z}}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)$ (13)
$\underset{\phantom{n\to\infty}}{=}\frac{|Z|}{n(n-1)}\left(\frac{1}{|Z|}\sum_{(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in Z}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)+\frac{|\bar{Z}|}{n(n-1)}\left(\frac{1}{|\bar{Z}|}\sum_{(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in\bar{Z}}S(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-})\right)$ (14)
=$\underset{n\to\infty}{=}\frac{1}{k}\cdot\mathbb{E}_{p(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-}\mid(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in Z)}[S]+\left(1-\frac{1}{k}\right)\cdot\mathbb{E}_{p(\hat{\mathbf{y}},\mathbf{y},\mathbf{y}_{-}\mid(\mathbf{y},\hat{\mathbf{y}},\mathbf{y}_{-})\in\bar{Z})}[S]$ (15)
@@
where $|Z|=n\left(n/k-1\right)$, and $|\bar{Z}|=n\left(n-n/k\right)$. We used the assumption of a large sample size at the third equality.
Recovery check of a single layer by iCNN
We performed a recovery check analysis using a single layer from the iCNN method in Fig 11. The iCNN method generates an image by optimizing pixel values to make the image’s latent features similar to the target latent features [29]. In the pixel optimization condition (the left columns of each recovery image in Fig 11), we directly optimized the pixel values of images to minimize the mean squared loss between the latent features of the image as well as the total-variance (TV) loss of pixel values [65].
Additionally, the iCNN method can incorporate image generator networks (middle and right columns of each recovery image in Fig 11) to add constraints on image statistics. Instead of optimizing the pixel values, we optimized the parameters related to the generator networks to minimize the mean squared loss between the latent features obtained through the generator networks and the target latent features. As a weak image prior, we used Deep Image Prior (DIP) [68]. DIP utilizes a hierarchical U-Net architecture as an inherent prior for image tasks, capturing the statistical regularities of images without relying on a specific dataset. This model works effectively by optimizing a randomly initialized neural network that can be used as an image prior in various inverse problems such as denoising, super-resolution, and inpainting tasks. In our analysis, DIP started with a U-Net initialized with random noise. Subsequently, the latent features and weight parameters of DIP were optimized to minimize the difference between the network’s output and target DNN features. For the pre-trained image prior [69], we used the same generator model as in Shen et al. [29] optimizing the latent features of the pre-trained networks.
Acknowledgments
@@
We thank our laboratory team, especially Eizaburo Doi, Haibao Wang, Kenya Otsuka, Hideki Izumi, and Matthias Mildenberger for their invaluable comments and suggestions on the manuscript. This work was supported by the Japan Society for the Promotion of Science (JSPS: KAKENHI grants JP20H05954, JP20H05705 to Y.K, JP21K17821 to Y.N, and 22KJ1801 to K.S.), Japan Science and Technology Agency (JST: CREST grants JPMJCR18A5, and JPMJCR22P3 to Y.K.), and New Energy and Industrial Technology Development Organization (NEDO: commissioned project, JPNP20006 to Y.K.).
Supporting information
S1 Fig. UMAP visualization of CLIP text features in the Natural Scene Dataset (NSD) using default parameters. Similar to Fig 4A in the main text, this scatter plot visualizes the distribution of semantic concepts within the NSD. The gray points represent training samples, and the orange points represent test samples. Unlike Fig 4A, which used parameters optimized for clustering visualization, this figure employs default UMAP settings. Despite the difference in parameters, the visualization still reveals distinct clusters and substantial overlap between training and test samples.
S2 Fig. UMAP visualization of CLIP text features in the Deeprecon dataset. This scatter plot visualizes the distribution of semantic concepts within the Deeprecon data, with the gray points representing training samples and the orange points indicating test samples. Unlike the NSD dataset (Fig 4A in the main text), this visualization demonstrates a clearer separation between training and test samples. This distinction highlights the Deeprecon dataset’s intentional design to differentiate object categories between training and test sets.
S3 Fig. Confusion matrices from the hold-out analyses. (A) Clustering results of CLIP text features. This panel shows the outcome of applying $k$-means clustering to the UMAP embedding of CLIP text features from the NSD dataset (Fig 4A in the main text). These clustering results were used to determine the assignment of training and test samples under the hold-out split condition (Fig 6A in the main text). (B) Confusion probability matrices for cluster identification. These matrices illustrate the performance of cluster identification under different conditions. The top row shows results under the hold-out split conditions, while the bottom row presents results under the naive split conditions. For each row, the left matrix represents CLIP text features, and the right matrix represents CLIP vision features. In each matrix, the cell at position $(i,j)$ indicates the proportion of samples originally from cluster $i$ that were predicted to belong to cluster $j$. As visualized in the confusion matrices, the hold-out split condition, where entire semantic clusters are excluded from training, reveals the challenges in zero-shot prediction of novel concepts. In contrast, the naive split condition, where samples from all clusters are included in the training, demonstrates the model’s performance when semantic concepts overlap between training and test sets.
S4 Fig. Extended simulation results for clustered data analysis. (A) Relationship between cluster identification accuracy and cluster variance ratio. Each subplot represents a different number of clusters used in model training. The $x$-axis represents the cluster variance ratio $\sigma_{\mathrm{intra}}^{2}/\sigma_{\mathrm{inter}}^{2}$, and the $y$-axis represents the cluster identification performance. The green lines represent results for in-distribution test samples, and the orange lines show outcomes for out-of-distribution (OOD) samples. As the cluster variance ratio increases, clusters tend to overlap more, leading to difficulty with identification in the in-distribution condition. However, larger cluster variance also allows for more extensive coverage of the feature space, potentially improving OOD prediction. Consequently, OOD curves typically show a peak at intermediate variance ratios, reflecting a balance between cluster distinctiveness and feature space coverage. (B) Heatmap of cluster identification accuracy. These matrices visualize how accuracy changes with varying numbers of training clusters and feature space dimensions. The cluster variance ratio ($\sigma_{\mathrm{intra}}^{2}/\sigma_{\mathrm{inter}}^{2}$) is fixed at $0.1$ for all simulations. The left matrix shows results for in-distribution samples, while the right matrix displays outcomes for out-of-distribution (OOD) samples. In each matrix, cell $(i,j)$ indicates the cluster identification accuracy when the feature dimension is $i$ and the number of clusters used in training is $j$. This view illustrates the interplay between feature dimensionality and the number of training clusters in determining the model’s generalization capabilities. Notably, in the OOD condition, as the feature dimensionality increases, a greater number of training clusters becomes necessary to achieve a robust zero-shot prediction.
@@
S5 Fig. Frame similarity analysis in the dataset of Nishimoto et al. [89]. Movie scenes were detected from the movie stimuli. Then, for the first and last frames of each test scene, we identified the frames from the training movies with the closest Euclidean distance. Our analysis revealed that $37$ out of the $48$ scenes in the test set contained frames that were nearly identical to those in the training set.